# Complete toric varieties with semisimple automorphism group

### v1

@@ 4.1. The case of the projective linear group
has cardinality $n+1-|\mathrm{supp}(\rho_{\alpha})|$ because:
@@
    When $k\in\mathrm{supp}(\rho_{\alpha})$, $\epsilon-a_{k}=0$, yielding identical roots $\epsilon e_{j}^{*}$
@@
    Distinct $k\notin\mathrm{supp}(\rho_{\alpha})$ produce distinct roots
Thus each set has size $n-|\mathrm{supp}(\rho_{\alpha})|+1$. By (5), the sets are disjoint across $j$, so:
$|\mathcal{R}(\rho_{\alpha})|=\sum_{j\in\mathrm{supp}(\rho_{\alpha})}(n+1-|\mathrm{supp}(\rho_{\alpha})|)=|\mathrm{supp}(\rho_{\alpha})|\cdot(n+1-|\mathrm{supp}(\rho_{\alpha})|).$
@@
$|\mathcal{R}(\rho)|=|\mathrm{supp}(\rho)|\cdot(n+1-|\mathrm{supp}(\rho)|)\geq n.$
Equality holds when $|\mathrm{supp}(\rho)|\in\{1,n\}$. In particular, $\rho$ has exactly $n$ associated roots if and only if $\rho=\pm e_{i}$ for some $i\in I_{n}^{0}$.
@@
Suppose $\Sigma$ is a fan with $\mathcal{R}(\Sigma)=R_{n}$. For $j\in I_{n}$, Proposition˜4.2 implies $\rho_{e_{j}^{*}}$ has the form $\sum_{k=1}^{n}a_{k}e_{k}$ with $a_{k}\in\{0,-1\}$ and $a_{j}=-1$. By Proposition˜4.5, its number of associated roots is:
$|\mathcal{R}(\rho_{e_{j}^{*}})|=|\mathrm{supp}(\rho_{e_{j}^{*}})|\cdot(n+1-|\mathrm{supp}(\rho_{e_{j}^{*}})|).$
@@
$|\mathcal{R}(\Sigma)|\geq n^{2}+|\mathrm{supp}(\rho_{e_{j}^{*}})|(n+1-|\mathrm{supp}(\rho_{e_{j}^{*}})|)>n^{2}+n=|\mathcal{R}_{n}|,$
@@
Suppose $\rho_{e_{j}^{*}}=-e_{j}$ for some $j$. Since $\langle e_{0},e_{j}^{*}\rangle=-1$, Definition˜3.1 forces $e_{0}\notin\Sigma(1)$. Thus $\rho_{e_{j}^{*}}=-e_{j}$ for all $j\in I_{n}$, giving $\{-e_{i}\mid i\in I_{n}\}\subseteq\Sigma(1)$. Note that $\rho_{\alpha}$ for $\alpha=e_{st}^{*}$ lies in this set, but $\rho_{-e_{i}^{*}}$ does not. Thus the only roots $e_{i}^{*}$ lack an associated ray in $\Sigma(1)$. Since each ray requires $\geq n$ roots, we need a ray lying in
@@ 4.2. The general case
We now consider semisimple $\operatorname{Aut}^{\circ}X$. By [Dem70, Section 3, §4, Corollart to Theorem 2], $\operatorname{Aut}^{\circ}_{X}$ decomposes as a product of projective linear groups. Combining results from the last subsection with combinatorial tools from Section 2, we establish propositions leading to Theorem˜1.2.
The proof strategy proceeds as follows. For each $i$, let $T_{i}$ be the standard maximal torus of $\operatorname{PGL}_{n_{i}+1}$, with character lattice $M_{i}$ and cocharacter lattice $N_{i}$. Then $X$ is a toric variety with:
@@
Let $\Sigma$ be the fan of $X$ in $V\coloneqq N\otimes_{\mathbb{Z}}\mathbb{R}=\bigoplus_{i=1}^{k}V_{i}$, and $R(\Sigma)\subseteq M$ its Demazure root set. Then $R(\Sigma)$ must coincide with the product of root data:
@@
We now derive consequences for $\Sigma_{X}$. Our objective is to prove that for each $i\in I_{k}$, the restricted fan $\Sigma|_{V_{i}}$ equals $\Sigma_{\mathbb{P}^{n_{i}}}$. This will follow from two subsequent lemmas.
Let $\rho\in\Sigma_{X}(1)$. Since $\rho\neq 0$, there exists some $t\in I_{k}$ such that $\pi_{t}(\rho)\neq 0$, where $\pi_{t}:N\to N_{t}$ is the projection. The following lemma explicitly describes $\pi_{t}(\rho)$.
@@
We conclude that if $n_{s}=n_{t}=1$, there cannot exist $\rho\in\Sigma_{X}(1)$ with both $\pi_{s}(\rho)\neq 0$ and $\pi_{t}(\rho)\neq 0$.
We now proceed to the case where $n_{s}>1$. Following the strategy of Subcase 2, we construct a root outside the root system at $X$. Observe that $\rho$ is associated with roots in the set:
@@
Subcase 2: $\pi_{s}(\rho^{\prime})=e_{0,s}$. In this case, we choose $\alpha_{s}=e_{1,s}^{*}$ and $\alpha_{t}=-e_{1,t}^{*}$. Then $\rho$, $\alpha_{t}$, $\rho^{\prime}$, and $\alpha_{s}$ satisfy condition (9).
Now assume $\pi_{t}(\rho^{\prime})\neq 0$. By Lemma˜4.9, we consider the following subcases:
@@
This proves we can always find $\rho^{\prime}$, $\alpha_{s}$, and $\alpha_{t}$ satisfying (9).
We now show $\alpha_{t}+\alpha_{s}$ is a Demazure root of $\Sigma_{X}$. First,

### v2

@@ 4.1. The case of the projective linear group
has cardinality $n+1-|\operatorname{supp}(\rho_{\alpha})|$ because:
@@
    When $k\in\operatorname{supp}(\rho_{\alpha})$, $\epsilon-a_{k}=0$, yielding identical roots $\epsilon e_{j}^{*}$
@@
    Distinct $k\notin\operatorname{supp}(\rho_{\alpha})$ produce distinct roots
Thus each set has size $n-|\operatorname{supp}(\rho_{\alpha})|+1$. By (5), the sets are disjoint across $j$, so:
$|\mathcal{R}(\rho_{\alpha})|=\sum_{j\in\operatorname{supp}(\rho_{\alpha})}(n+1-|\operatorname{supp}(\rho_{\alpha})|)=|\operatorname{supp}(\rho_{\alpha})|\cdot(n+1-|\operatorname{supp}(\rho_{\alpha})|).$
@@
$|\mathcal{R}(\rho)|=|\operatorname{supp}(\rho)|\cdot(n+1-|\operatorname{supp}(\rho)|)\geq n.$
Equality holds when $|\operatorname{supp}(\rho)|\in\{1,n\}$. In particular, $\rho$ has exactly $n$ associated roots if and only if $\rho=\pm e_{i}$ for some $i\in I_{n}^{0}$.
@@
Suppose $\Sigma$ is a fan with $\mathcal{R}(\Sigma)=\mathcal{R}_{n}$. For $j\in I_{n}$, Proposition˜4.2 implies $\rho_{e_{j}^{*}}$ has the form $\sum_{k=1}^{n}a_{k}e_{k}$ with $a_{k}\in\{0,-1\}$ and $a_{j}=-1$. By Proposition˜4.5, its number of associated roots is:
$|\mathcal{R}(\rho_{e_{j}^{*}})|=|\operatorname{supp}(\rho_{e_{j}^{*}})|\cdot(n+1-|\operatorname{supp}(\rho_{e_{j}^{*}})|).$
@@
$|\mathcal{R}(\Sigma)|\geq n^{2}+|\operatorname{supp}(\rho_{e_{j}^{*}})|(n+1-|\operatorname{supp}(\rho_{e_{j}^{*}})|)>n^{2}+n=|\mathcal{R}_{n}|,$
@@
Suppose $\rho_{e_{j}^{*}}=-e_{j}$ for some $j$. Since $\langle e_{0},e_{j}^{*}\rangle=-1$, Definition˜3.1 forces $e_{0}\notin\Sigma(1)$. Thus $\rho_{e_{j}^{*}}=-e_{j}$ for all $j\in I_{n}$, giving $\{-e_{i}\mid i\in I_{n}\}\subseteq\Sigma(1)$. Note that $\rho_{\alpha}$ for $\alpha=e_{st}^{*}$ lies in this set, but $\rho_{-e_{i}^{*}}$ does not. Thus the only roots $e_{i}^{*}$ are the only roots that lack an associated ray. Since each ray requires $\geq n$ roots, we need a ray lying in
@@ 4.2. The general case
We now consider semisimple $\operatorname{Aut}^{\circ}X$. By [Dem70, Section 3, §4, Corollary to Theorem 2], $\operatorname{Aut}^{\circ}_{X}$ decomposes as a product of projective linear groups. Combining results from the last subsection with combinatorial tools from Section 2, we establish propositions leading to Theorem˜1.2. The proof strategy proceeds as follows. For each $i$, let $T_{i}$ be the standard maximal torus of $\operatorname{PGL}_{n_{i}+1}$, with character lattice $M_{i}$ and cocharacter lattice $N_{i}$. Then $X$ is a toric variety with:
@@
Let $\Sigma$ be the fan of $X$ in $V\coloneqq N\otimes_{\mathbb{Z}}\mathbb{R}=\bigoplus_{i=1}^{k}V_{i}$, and $\mathcal{R}(\Sigma)\subseteq M$ its Demazure root set. Then $\mathcal{R}(\Sigma)$ must coincide with the product of root data:
@@
We now derive consequences for $\Sigma_{X}$. Our objective is to prove that for each $i\in I_{k}$, the restricted fan $\Sigma|_{V_{i}}$ equals $\Sigma_{\mathbb{P}^{n_{i}}}$. This will follow from two subsequent lemmas. Let $\rho\in\Sigma_{X}(1)$. Since $\rho\neq 0$, there exists some $t\in I_{k}$ such that $\pi_{t}(\rho)\neq 0$, where $\pi_{t}:N\to N_{t}$ is the projection. The following lemma explicitly describes $\pi_{t}(\rho)$.
@@
We conclude that if $n_{s}=n_{t}=1$, there cannot exist $\rho\in\Sigma_{X}(1)$ with both $\pi_{s}(\rho)\neq 0$ and $\pi_{t}(\rho)\neq 0$. We now proceed to the case where $n_{s}>1$. Following the strategy of Subcase 2, we construct a root outside the root system at $X$. Observe that $\rho$ is associated with roots in the set:
@@
Subcase 2: $\pi_{s}(\rho^{\prime})=e_{0,s}$. In this case, we choose $\alpha_{s}=e_{1,s}^{*}$ and $\alpha_{t}=-e_{1,t}^{*}$. Then $\rho$, $\alpha_{t}$, $\rho^{\prime}$, and $\alpha_{s}$ satisfy condition (9). Now assume $\pi_{t}(\rho^{\prime})\neq 0$. By Lemma˜4.9, we consider the following subcases:
@@
This proves we can always find $\rho^{\prime}$, $\alpha_{s}$, and $\alpha_{t}$ satisfying (9). We now show $\alpha_{t}+\alpha_{s}$ is a Demazure root of $\Sigma_{X}$. First,